If the regression line y ̂ = 3 + 2x has been fitted to the data points (4, 8), (2, 5), and (1, 6), the residual sum of squares will be:  ?
Answer #1

Residual = y - 3 - 2x

squared residual = (y-3-2x)^2

for (4,8) it is 9

for (2,5) it is 4

for (1,6) it is 1

sum is 9+4+1 = 14

The residual sum of square =14